### v1

# Towards deterministic non-Gaussianity on a chip

Abstract
We propose cascaded spontaneous four-wave mixing (SFWM) in microring resonators as a scalable and efficient approach for directly generating non-Gaussian states of light. Focusing on the well-understood “low-gain” regime, we demonstrate that triplet generation through cascaded SFWM can be achieved with high efficiency and favorable spectral characteristics using realistic microring sources in AlGaAs. The ability to achieve the generation of light in a single set of supermodes – and the accessibility of the “high-gain” regime at realistic pump powers – makes this source a promising candidate as a direct and deterministic source of non-Gaussian light for photonic quantum information processing.
preprint: APS/123-QED
Introduction—Generating non-Gaussian states of light is a central topic in the quantum optics and photonic quantum information processing communities. Much of the work has been motivated by underlying fundamental physics: Due to their uniquely “non-classical” characteristics, non-Gaussian states are a necessary resource for implementing photonic quantum information processing [1]. But creating non-Gaussian states is challenging [2]. The conventional approach is to apply measurements – usually photon-number-resolving (PNR) detection – to squeezed light [3, 4, 5]. Although PNR detectors and squeezed light sources are readily available, relying on detection to introduce non-Gaussianity has the drawbacks of being probabilistic, introducing latency times, and requiring cryogenic components in a setting that could otherwise operate at room temperature. This has led to a search for deterministic, room-temperature sources of non-Gaussian light based on the use of parametric processes [6].
Any such source must satisfy at least three requirements to be of broad practical interest. First, it must have a sufficiently high brightness. Second, for the generated light to be useful and easily characterized it must be described by a small set of supermodes [7]. Finally, scalability – for example, through on-chip integration – is required for the source to be relevant in the large-scale information-processing applications that require non-Gaussianity as a resource.
For a parametric nonlinearity to induce non-Gaussianity, it must be described by a Hamiltonian that is at least cubic in bosonic operators [8]. One set of proposals involves the use of third-order parametric down-conversion (TOPDC), a non-Gaussian unitary process that can be described as a higher-order squeezing interaction in which photons are produced in sets of triplets. Many TOPDC-based sources have been proposed, yet this approach is hampered by the challenge of achieving phase matching, and the weakness of the $\chi^{(3)}$ nonlinearity on which the process relies.
Another proposal relies on pump depletion and the resulting entanglement of pump light with photon pairs generated by the pump. This has been theoretically considered in the so-called mesoscopic regime in scenarios involving long waveguides [9, 10], and in microring resonators requiring both very high finesse and pump powers [11]. Experimental implementation of these schemes have been hampered by fabrication requirements and material limitations, and issues of scalability and few-supermode generation can also be of concern.
An alternative strategy would be the use of cascaded squeezing, where a squeezed vacuum is generated and then used to drive another squeezed light source; because the second process is driven by a non-classical pump, it has the desired non-Gaussian form [9, 6]. If the first squeezing interaction is restricted to the low-gain regime, this results in the probabilistic generation of photon triplets: This process – and the certification of the resulting tripartite entanglement – has been implemented through spontaneous parametric down-conversion (SPDC) in bulk crystals [12, 13, 14, 15] and in integrated waveguides [16]. Besides the inherent scalability, the use of integrated structures offers an enhanced effective nonlinearity due to the strong transverse field confinement.
Yet implementations of cascaded squeezing in microring resonators have not yet been investigated in detail, despite the fact that microring sources satisfy the three requirements mentioned above. First, they provide an additional resonant field enhancement, further increasing the structure’s effectiveness in light generation [17]. We will also show that the type of “few-supermode” operation we seek can be obtained very effectively using resonant sources, often achieving a higher level of separability than non-resonant sources [7, 18]. Finally, the scalability of microring resonators has already made them a popular choice in the implementation of large-scale integrated devices.
The enhanced nonlinearity in these devices allows one to envision implementing cascaded spontaneous four-wave mixing (SFWM) rather than SPDC. This has a number of practical advantages: The energy conservation conditions in SFWM allow for all the fields to be very close to each other in frequency, for instance, in the same telecom band. Phase matching – which can be a significant challenge in triplet generation – is therefore widely accessible. The narrow frequency range should also make it relatively easy to find appropriate material platforms that are transparent at the relevant wavelengths and that do not exhibit deleterious effects such as two-photon absorption.
For the sample calculations, we envision the use of integrated AlGaAs structures, but the scheme we propose can be implemented in other platforms. As well, to benchmark the results that can be expected in first experiments, we focus on the probabilistic generation of photon triplets, although with reasonable experimental parameters one can go beyond that limit.
Cascaded spontaneous four-wave mixing—We propose a source consisting of two rings (see Fig. 1a), such that the first (second) pair generation process occurs in the first (second) ring: This is ensured by tuning the positions of the ring resonances, resulting in a flexible and practical design in which each ring’s resonance frequencies, free spectral range (FSR), and waveguide coupling can be controlled independently. Although single-ring cascaded SFWM could be implemented in principle, this would be challenging, and it would require careful dispersion engineering to ensure the proper resonance alignment. A device with two rings also allows for the two pump fields to be confined to separate rings, preventing some of the parasitic processes that would occur if both interactions occurred in a single ring. The basic microring structures sketched in Fig. 1a are sufficient for demonstrations of cascaded SFWM, but in principle they could be replaced by more sophisticated structures – e.g., with interferometric couplers or photonic molecules – to provide even more tunability of individual resonances, and suppression of unwanted nonlinear processes [19, 20, 21, 22].
We envision working at telecom wavelengths, and we focus on a scheme in which the generated modes are non-degenerate in frequency (see Fig. 2a). We choose this configuration to clearly identify the different modes involved in the interactions, but the design can be modified to affect the positions of the pump and generated fields; for example, degenerate triplet sources can be designed by adjusting the positions of the rings’ resonances (see Fig. 2b).
Figure 1: a) Subsequent microring resonators coupled to a bus waveguide, each coupled to a phantom waveguide to model photon losses. b) Electric field ($E_{x}$) mode profile of the TE00 mode. Mode simulations are done in Lumerical with a bend radius of 10 $\mu\mathrm{m}$, using the refractive index for Al0.3Ga0.7As from [23].
Figure 2: Energy diagrams for the generation of non-degenerate triplets (a), and degenerate triplets (b). The SFWM process for the first (second) ring is given in each subfigure’s top (bottom) panel.
The Hamiltonian describing the nonlinearity is given by [24]
$H_{\mathrm{NL}}=-\frac{1}{4\epsilon_{0}}\int\mathrm{d}\bm{r}\Gamma^{(3)}_{ijkl}(\bm{r})D^{i}(\bm{r})D^{j}(\bm{r})D^{k}(\bm{r})D^{l}(\bm{r}),$ (1)
where repeated Cartesian indices are implicitly summed over and where $\bm{D}(\bm{r})$ denotes the displacement field in the structure [24], which we expand in terms of asymptotic fields [25, 26] having the form
$\bm{D}_{u}^{\mathrm{in}}(\bm{r})=\sum_{\lambda}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\mathrm{in},\lambda}(\bm{r})a_{u}^{\mathrm{in},\lambda}(k)+\mathrm{H.c.,}$ (2)
and similarly for the “out” fields, and we show the linear Hamiltonian in Appendix Towards deterministic non-Gaussianity on a chip. The index $\lambda$ denotes an input or output channel through which light can propagate [27, 28], and for the structure given in Fig. 1a, $\lambda\in\{\mathrm{ac,ph(1),ph(2)}\}$ for the actual, first, and second phantom channels; the phantom channels are used to account for scattering losses [25]. The $a_{u}^{\text{in/out},\lambda}(k)$ denote bosonic lowering operators with the usual commutation relations [27], and $\bm{\mathcal{D}}_{uk}^{\text{in/out},\lambda}(\bm{r})$ denotes the asymptotic field amplitude. The form of the latter depends on the structure in question (see Supplemental Material 1).
We restrict ourselves to the terms in the full nonlinear Hamiltonian relevant to cascaded SFWM: $H_{\mathrm{NL}}^{(I)}(t)=H_{\mathrm{SFWM1}}^{(I)}(t)+H_{\mathrm{SFWM2}}^{(I)}(t)$ (see Appendix Towards deterministic non-Gaussianity on a chip). The state generated by the nonlinear interactions is given by
[…]
$+\sigma\sum_{\mu\mu^{\prime}\mu^{\prime\prime}}\int\mathrm{d}k_{1}\mathrm{d}k_{2}\mathrm{d}k_{3}\Psi^{\mu\mu^{\prime}\mu^{\prime\prime}}(k_{1},k_{2},k_{3};t\rightarrow\infty)$
[…]
where we have dropped the “out” label from the operators. The second term in Eq. (4) refers to the generation of a photon pair in the first SFWM process. As described earlier [27], $\varphi^{\lambda\lambda^{\prime}}(k_{1},k_{2};t)$ denotes the biphoton wavefunction (BWF) that describes the correlations between a pair of photons that couple out of the device via the channels $\lambda$ and $\lambda^{\prime}$, and is given in Eq. (31).
The final term in Eq. (4) refers to the generation of a photon triplet through cascaded SFWM. The spectral correlations for a photon triplet are characterized by the triphoton wavefunction (TWF), given in Eq. (32).
The probabilities per pump pulse of generating pairs from the first process and triplets from the cascaded processes are $|\beta|^{2}$ and $|\sigma|^{2}$, respectively: These are computed by imposing the normalization conditions in Eqs. (33)–(34) on the BWF and TWF. A corresponding probability for photon pairs coupling out through channels $\lambda$ and $\lambda^{\prime}$ can be denoted $|\beta^{\lambda\lambda^{\prime}}|^{2}$; these quantities can be obtained from the BWF as described earlier [28]. Similarly, one can also compute specific probabilities $|\sigma^{\mu\mu^{\prime}\mu^{\prime\prime}}|^{2}$ to evaluate, for example, the rate of threefold coincidences in the physical output channel. The state of interest for triphoton detection is the one where all three photons are at the physical output channel (i.e. $\mu=\mu^{\prime}=\mu^{\prime\prime}=\mathrm{ac}$), and is characterized by $\Psi^{\mathrm{ac,ac,ac}}(k_{1},k_{2},k_{3};t\rightarrow\infty)$. This state can be written as
[…]
Few-supermode operation—The ability to generate light in the minimal number of supermodes is essential in developing a useful source of non-Gaussianity, just as one seeks single-Schmidt-mode sources of squeezed light [29, 7, 30, 31, 32]. In the low-gain regime this is achieved by obtaining a TWF that is separable [7], and although more work is needed to understand non-Gaussian nonlinear optics in the high gain regime, this approach can be expected to be useful there as well.
The separability of the TWF can be characterized by referring to the purities of the reduced density matrices
$\rho^{(1)}(k_{1},k_{1}^{\prime})$ $=|\mathcal{N}|^{2}\int\mathrm{d}k_{2}\mathrm{d}k_{3}\Psi^{\mathrm{ac,ac,ac}}(k_{1},k_{2},k_{3};t\rightarrow\infty)$
$\hskip 34.14322pt\times\left(\Psi^{\mathrm{ac,ac,ac}}(k_{1}^{\prime},k_{2},k_{3};t\rightarrow\infty)\right)^{*}$ (6)
(and similarly for $\rho^{(2)}$ and $\rho^{(3)}$) [7]. In the limit of a perfectly separable TWF, implying the light is generated in a single set of supermodes, $p_{1}\equiv\mathrm{Tr}\left(\rho^{(1)}(k_{1},k_{1}^{\prime})\right)=1$. Thus, to design a source with our desired mode structure, we seek a configuration in which $p_{1}$, $p_{2}$, and $p_{3}$ are as close as possible to unity. In other contexts, it is known that one can shape the joint spectral amplitude produced by a resonant device by manipulating the pump durations [33, 28]; in the following sample calculations, we show that this strategy also applies to the TWF in cascaded SFWM.
Sample calculation—We now compute triplet generation TWFs and rates for a realistic device. We consider implementations in the AlGaAs platform: This emerging platform supports large nonlinear coefficients $\gamma_{\mathrm{NL}}$, has demonstrated high quality factors [34], and is compatible with hybrid integration approaches [35]. In our calculations, we take $\chi^{(3)}\approx 5.6\times 10^{-19}$ m2/V2 ($n_{2}\approx 1.43\times 10^{-17}$ m2/W) [36, 34, 35], corresponding to nonlinear parameters of $\gamma_{\mathrm{NL}}\approx 250\,(\mathrm{W\cdot m})^{-1}$ for both SFWM processes, computed using the TE00 modes (see Fig. 1b and Supplemental Material 2). For the ring dimensions we take $R_{1}={20\,\mu\mathrm{m}}$ and $R_{2}={10\,\mu\mathrm{m}}$ (recall the labelling in Fig. 1a), both with cross-sectional dimensions of $(400\times 920)$ nm (see Fig. 1b). We assume realistic intrinsic quality factors of $Q_{\mathrm{int}}=10^{6}$ for all the resonances in the frequency range of interest [34]. The rings and pumps should be tuned such that both pumps are on resonance.
We consider the pump fields $P_{1}$ and $P_{2}$ for the first and second ring respectively (recall Fig. 1 & 2), to be centered at wavelengths $\lambda_{P_{1}}=1546.83$ nm and $\lambda_{P_{2}}=1584.29$} nm, respectively. The triplets are generated in the resonances with center wavelengths $\lambda_{S_{1}}=1541.60$ nm, $\lambda_{S_{2}}=1562.68$ nm, and $\lambda_{S_{3}}=1573.41$ nm. The rate of triplets at the chip’s output is given by $R_{3}=R_{R}\eta_{S_{1}}^{\mathrm{ac}(1)}\eta_{S_{2}}^{\mathrm{ac}(2)}\eta_{S_{3}}^{\mathrm{ac}(2)}|\sigma|^{2}$, where $R_{R}$ is the pulses’ repetition rate, and $\eta_{u}^{\lambda(i)}$ are the escape efficiencies (see Eq. (17)).
To ensure the validity of our perturbative treatment (recall Eq. (4)), we limit the average power in the first pump $P_{1}$ to yield a pair generation probability of $|\beta|^{2}\approx 0.1$. The power in the second pump $P_{2}$ will be set to avoid any material damage due to linear and nonlinear absorption [37, 38, 39]. The pump durations and relative escape efficiencies impact the triphoton wavefunctions and rates; we show this by considering three different pumping configurations, summarized in Tab. 1.
Configurations of first pump ($P_{1}$) / second pump ($P_{2}$)
Configuration (AA) ($B$) ($C$)
Duration (FWHM) 300 ps / CW 300 ps / 300 ps 50 ps / 100 ps
Pulse energy 0.17 pJ / NA 0.17 pJ / 36.7 pJ 0.52 pJ / 20.5 pJ
Repetition rate 10 MHz / NA 10 MHz / 10 MHz 10 MHz / 10 MHz
Peak power in waveguide 0.53 mW / 1.4 mW 0.53 mW / 0.12 W 9.74 mW / 0.19 W
[…]
Purity $p_{1}$ / $p_{2}$ / $p_{3}$ 0.974 / 0.610 / 0.610 0.982 / 0.806 / 0.806 0.995 / 0.970 / 0.970
[…]
Rate 7.43 Hz 19.2 Hz 5.66 Hz
Table 1: Simulation results for different pump configurations, with $Q_{\mathrm{int}}=10^{6}$. Critical coupling ($\eta=0.5$) is used for configurations $A$ and $B$, and $\eta_{P_{1}}^{\mathrm{ac}(1)}=0.95$, $\eta_{S_{1}}^{\mathrm{ac}(1)}=\eta_{I}^{\mathrm{ac}(1)}=0.9$, $\eta_{P_{2}}^{\mathrm{ac}(2)}=\eta_{I}^{\mathrm{ac}(2)}=0.85$, and $\eta_{S_{2}}^{\mathrm{ac}(2)}=\eta_{S_{3}}^{\mathrm{ac}(2)}=0.5$ is used for configuration $C$.
Initially we consider two configurations $A$ and $B$, with critical coupling for all resonances, such that all loaded quality factors are $5\times 10^{5}$. We take $P_{1}$ to be a Gaussian pulse with full-width-half-maximum (FWHM) of $300$ ps – similar to the ring’s dwelling time – with a $R_{R}=10$ MHz repetition rate. Interesting triplet rates can be obtained with either pulsed or CW excitation in the second ring. In configuration $A$, $P_{2}$ is CW; in configuration $B$, it is a train of $300$ ps pulses. To obtain $|\beta|^{2}\approx 0.1$, we limit $P_{1}$ to a peak power of $0.53\,\mathrm{mW}$ in the bus waveguide. In configuration $A$, we take the power of the CW $P_{2}$ to be 1.4{1.4} mW, which corresponds to $\sim 1$ W circulating in the ring; this results in an output triplet rate of $R_{3}\approx{7.43}$ Hz. In configuration $B$, the pulsed $P_{2}$ is set such that the peak power in the ring does not exceed 10 W; this results in a peak power in the waveguide of 0.12 W, giving $R_{3}=19.2$ Hz.
We represent the TWFs obtained with configurations $A$ and $B$ in Fig. 3 (a) and (b), respectively. In moving from configuration $A$ to configuration $B$, we see that – similar to photon-pair generation in a ring resonator [40] – pumping with a pulse duration shorter than the dwelling time in the ring results in a TWF with reduced spectral correlations. The visualizations of the TWFs in Fig. 3 are isosurfaces at 10% of the modulus squared of the maximal value of the given TWF, and projections over each variable are provided on the panels.
[…]
Figure 3: Visualization of the total triphoton wavefunction $|\Psi|^{2}\equiv|\Psi^{\mathrm{ac,ac,ac}}(k_{1},k_{2},k_{3};t\rightarrow\infty)|^{2}$ for configuration AA (a), $B$ (b) and $C$ (c); we use $x_{i}=v_{S_{i}}(k_{i}-K_{S_{i}})/\bar{\Gamma}_{S_{i}}$, where $K_{S_{i}}$ indicates the reference wavenumber for the $i^{th}$ signal, and in the panel perpendicular to the $x_{i}$ axis the quantity indicated above the color bar is plotted using the scale of the bar.
Discussion—Our proposed device demonstrates that photon triplet generation can be achieved on a platform that is simultaneously efficient, compact, and readily manufacturable using existing fabrication techniques. Our proposed design requires lower average pump powers compared to other proposals, requiring only average on-chip powers of $\sim$ 1 mW or less, and we have used well-accepted material parameters. Furthermore, our resonant photon-triplet source offers both spectral tunability and control over the number of supermodes, providing essential flexibility for tailoring the properties of the generated non-Gaussian states. Our proposed device could obtain triplet rates comparable to those quoted in other theoretical proposals [42, 43, 44, 45], while being more efficient, compact, and allowing for tunability of the light’s spectral properties. For comparison to experimentally demonstrated sources, taking a loss of 3 dB per output channel [46, 47] for a total loss of 9 dB; this would result in a measured triplet coincidence rate of $\sim 1$ Hz, higher than the rates observed in experimental implementations of cascaded pair generation [12, 13, 16, 14]. In addition, we note that the proposed device could achieve higher triplet rates by further increasing the repetition rate, since the resonator’s dwelling times (maximally $\sim 400$ ps for critical coupling) do not limit the repetition rate to the 10 MHz used in our calculations. We believe these high rates and promising characteristics strongly motivate the fabrication and demonstration of these sources, including the characterization of the generated light, which could be done through stimulated emission tomography [48], for example.
Perhaps most interesting is the fact that this source, due to its high efficiency, is a strong candidate for implementing cascaded SFWM in a high-gain regime: Above, we restricted the first SFWM process to a low-gain regime by restricting the energy of $P_{1}$ to less than a picojoule. From an experimental perspective, one can easily envision increasing the strength of $P_{1}$, resulting in squeezed vacuum which would then drive the second SFWM process, with the possibility of measurable non-Gaussianity in the field generated from the latter process. The dynamics of cascaded SFWM in this high-gain regime, and the characteristics of the deterministically generated non-Gaussian state it would produce, will be a rich and challenging topic for future theoretical work. It will also be a novel source of non-classical light in more complex quantum photonic devices [6, 49, 50].
Acknowledgements.
S.E.F. and J.E.S. acknowledge the Natural Sciences and Engineering Research Council of Canada European Union’s Horizon Europe Research and Innovation Programme (101070700, project MIRAQLS) for financial support, and thank Colin Vendromin for insightful discussions. S.E.F. acknowledges support from a Walter C. Sumner Memorial Fellowship. M.L. acknowledges support by PNRR MUR Project No. PE0000023-NQSTI. M.B. acknowledges support from the Quantum Research and Development Initiative, led by the National Research Council Canada, under the National Quantum Strategy, and thanks Nicolás Quesada and Boris Braverman for stimulating discussions.
Appendix B: Nonlinear physics— The relevant Hamiltonians for the cascaded SFWM processes are
$H_{\text{SFWM1}}$ $=-\frac{1}{4\epsilon_{0}}\frac{4!}{2!1!1!}\int\mathrm{d}\bm{r}\Gamma^{(3)}_{ijkl}(\bm{r})D^{\text{out},i}_{S_{1}}(\bm{r})D^{\text{out},j}_{I}(\bm{r})$
$\hskip 85.35826pt\times D^{\text{in},k}_{P_{1}}(\bm{r})D^{\text{in},l}_{P_{1}}(\bm{r}),$ (18)
$H_{\text{SFWM2}}$ $=-\frac{1}{4\epsilon_{0}}\frac{4!}{1!1!1!1!}\int\mathrm{d}\bm{r}\Gamma^{(3)}_{ijkl}(\bm{r})D^{\text{out},i}_{S_{2}}(\bm{r})D^{\text{out},j}_{S_{3}}(\bm{r})$
$\hskip 85.35826pt\times D^{\text{out},k}_{I}(\bm{r})D^{\text{in},l}_{P_{2}}(\bm{r}).$ (19)
We move to the interaction picture using $H_{\mathrm{NL}}^{(I)}(t)=e^{iH_{\mathrm{L}}t/\hbar}H_{\mathrm{NL}}e^{-iH_{\mathrm{L}}t/\hbar}$, where $H_{\mathrm{L}}$ is the linear Hamiltonian. In terms of the asymptotic field operators, we can write the linear Hamiltonian as
$H_{\mathrm{L}}=\sum_{u,\lambda}\int\mathrm{d}k\hbar\omega_{uk}\left(a_{u}^{\text{in/out},\lambda}(k)\right)^{\dagger}a^{\text{in/out},\lambda}_{u}(k),$ (20)
so the time dependence of the ladder operators in the interaction picture is trivial; we have
$\bm{D}_{u}^{\text{in/out}(I)}(\bm{r},t)=$ $\sum_{\lambda}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\text{in/out},\lambda}(\bm{r})$
$\times a_{u}^{\text{in/out},\lambda}(k)e^{-i\omega_{uk}t}+\mathrm{H.c..}$ (21)
We treat the pump fields $P_{1}$ and $P_{2}$ classically, taking $a_{u}^{\mathrm{in,ac}}(k)\rightarrow\alpha_{u}\phi_{u}(k)$, where $\phi_{u}(k)$ is a normalized function that defines the pump spectrum, and $\alpha_{u}$ is a constant set to yield the correct pulse energy. In this work, we only have pump fields entering through one channel (the bus waveguide), so the asymptotic-in fields for the pump fields have the form
$\bm{D}_{u}^{\text{in}(I)}(\bm{r},t)=\alpha_{u}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\text{in},\lambda}(\bm{r})\phi_{u}(k)e^{-i\omega_{uk}t}+\mathrm{H.c..}$ (22)
With the fields in the interaction picture, and recalling Eq. (18), we find
$H_{\text{SFWM1}}^{(I)}$ $(t)=\frac{-\alpha_{P_{1}}^{2}}{3\epsilon_{0}}\int\mathrm{d}\bm{r}\Gamma_{ijkl}^{(3)}(\bm{r})$
$\times$ $\left(\sum_{\lambda}\int\mathrm{d}k{\mathcal{D}}_{S_{1}k}^{\text{out},\lambda,i}(\bm{r})a_{S_{1}}^{\text{out},\lambda}(k)e^{-i\omega_{S_{1}k}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\sum_{\lambda^{\prime}}\int\mathrm{d}k{\mathcal{D}}_{Ik}^{\text{out},\lambda^{\prime},j}(\bm{r})a_{I}^{\text{out},\lambda^{\prime}}(k)e^{-i\omega_{Ik}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\int\mathrm{d}k{\mathcal{D}}_{P_{1}k}^{\text{in},k}(\bm{r})\phi_{P_{1}}^{\text{in}}(k)e^{-i\omega_{P_{1}k}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\int\mathrm{d}k{\mathcal{D}}_{P_{1}k}^{\text{in},l}(\bm{r})\phi_{P_{1}}^{\text{in}}(k)e^{-i\omega_{P_{1}k}t}+\mathrm{H.c.}\right).$ (23)
Rearranging and neglecting the fast-rotating terms, we have
$H^{(I)}_{\text{SFWM1}}(t)$ $=-\sum_{\lambda\lambda^{\prime}}\int\mathrm{d}k_{1}\mathrm{d}k_{2}\mathrm{d}k_{3}\mathrm{d}k_{4}K^{\lambda\lambda^{\prime}}_{1}(k_{1},k_{2},k_{3},k_{4})$
$\times e^{-i\Omega_{1}(k_{1},k_{2},k_{3},k_{4})t}\alpha^{2}_{P_{1}}\phi_{P_{1}}(k_{3})\phi_{P_{1}}(k_{4})$
$\times\left[a^{\text{out},\lambda}_{S_{1}}(k_{1})a^{\text{out},\lambda^{\prime}}_{I}(k_{2})\right]^{\dagger},$ (24)
and similarly, using Eq. (19), for the second SFWM interaction, we obtain
$H^{(I)}_{\text{SFWM2}}(t)$ $=-\sum_{\mu\mu^{\prime}\mu^{\prime\prime}}\int\mathrm{d}k_{1}\mathrm{d}k_{2}\mathrm{d}k_{3}\mathrm{d}k_{4}K^{\mu\mu^{\prime}\mu^{\prime\prime}}_{2}(k_{1},k_{2},k_{3},k_{4})$
$\times e^{-i\Omega_{2}(k_{1},k_{2},k_{3},k_{4})t}\alpha_{P_{2}}\phi_{P_{2}}(k_{4})$
$\times a^{\text{out},\mu}_{I}(k_{3})\left[a^{\text{out},\mu^{\prime}}_{S_{2}}(k_{1})a^{\text{out},\mu^{\prime\prime}}_{S_{3}}(k_{2})\right]^{\dagger}.$ (25)
We have introduced
$\Omega_{1}(k_{1},...,k_{4})$ $=\omega_{P_{1}k_{3}}+\omega_{P_{1}k_{4}}-\omega_{S_{1}k_{1}}-\omega_{Ik_{2}}$ (26)
$\Omega_{2}(k_{1},...,k_{4})$ $=\omega_{Ik_{3}}+\omega_{P_{2}k_{4}}-\omega_{S_{2}k_{1}}-\omega_{S_{3}k_{2}}$ (27)
and
$K_{1}^{\lambda\lambda^{\prime}}(k_{1},...,k_{4})=\frac{3}{\epsilon_{0}}\int\mathrm{d}\bm{r}$ $\Gamma_{ijkl}^{(3)}\left[\mathcal{D}^{\text{out},\lambda,i}_{S_{1}k_{1}}\mathcal{D}^{\text{out},\lambda^{\prime},j}_{Ik_{2}}\right]^{*}$
$\times\mathcal{D}^{\text{in},k}_{P_{1}k_{3}}\mathcal{D}^{\text{in},l}_{P_{1}k_{4}}$ (28)
$K_{2}^{\mu\mu^{\prime}\mu^{\prime\prime}}(k_{1},...,k_{4})=\frac{6}{\epsilon_{0}}\int\mathrm{d}\bm{r}$ $\Gamma_{ijkl}^{(3)}\left[\mathcal{D}^{\text{out},\mu^{\prime},i}_{S_{2}k_{1}}\mathcal{D}^{\text{out},\mu^{\prime\prime},j}_{S_{3}k_{2}}\right]^{*}$
$\times\mathcal{D}^{\text{out},\mu,k}_{Ik_{3}}\mathcal{D}^{\text{in,ac},l}_{P_{2}k_{4}},$ (29)
where the position dependence of the integrands was dropped for brevity.
Focusing on the “low-gain” regime for both SFWM processes, from Eq. (3) the generated state can be written as
[…]
We neglect the final term in Eq. (30), which describes the generation of two pairs in the first ring. Inserting Eqs. (24) and (25) into (30) and taking $_{0}\rightarrow-\infty, tt\rightarrow\infty$, we obtain Eq. (4), with
$\varphi^{\lambda\lambda^{\prime}}(k_{1},k_{2};t)$ $=\frac{i\alpha_{P_{1}}^{2}}{\hbar\beta}\int\mathrm{d}k_{3}\mathrm{d}k_{4}K_{1}^{\lambda\lambda^{\prime}}(k_{1},...,k_{4})\phi_{P_{1}}(k_{3})$
$\times\phi_{P_{1}}(k_{4})\int_{-\infty}^{t}\mathrm{d}t^{\prime}e^{-i\Omega_{1}(k_{1},...,k_{4})t^{\prime}}$ (31)
and
$\Psi^{\mu\mu^{\prime}\mu^{\prime\prime}}(k^{\prime}_{1},k_{2},k_{3};t)$ $=-\frac{i\alpha_{P_{2}}\beta}{\hbar\sigma}\int\mathrm{d}k_{1}\mathrm{d}k_{4}K_{2}^{\mu\mu^{\prime}\mu^{\prime\prime}}(k_{1},...,k_{4})$
$\times\phi_{P_{2}}(k_{4})\int_{-\infty}^{t}\mathrm{d}t^{\prime}\sum_{\lambda^{\prime}}\varphi^{\mu\lambda^{\prime}}(k^{\prime}_{1},k_{1};t^{\prime})$
$\times e^{-i\Omega_{2}(k_{1},...,k_{4})t^{\prime}}.$ (32)
The wavefunctions are normalized according to
$\sum_{\lambda\lambda^{\prime}}\int\mathrm{d}k_{1}\mathrm{d}k_{2}|\varphi^{\lambda\lambda^{\prime}}(k_{1},k_{2};t\rightarrow\infty)|^{2}=1$ (33)
and
$\hskip-8.5359pt\sum_{\mu\mu^{\prime}\mu^{\prime\prime}}\int\mathrm{d}k_{1}\mathrm{d}k_{2}\mathrm{d}k_{3}|\Psi^{\mu\mu^{\prime}\mu^{\prime\prime}}(k_{1},k_{2},k_{3};t\rightarrow\infty)|^{2}=1.$ (34)
Appendix C Supplemental material 1: Asymptotic fields for dual rings coupled to a waveguide
The asymptotic-in and -out mode fields we consider form a complete set for the structures of interest, and we can write the total displacement field
$\bm{D}(\bm{r})=\sum_{u}\bm{D}_{u}^{\mathrm{in/out}}(\bm{r})+\mathrm{H.c.},$ (35)
with
$\bm{D}_{u}^{\mathrm{in/out}}(\bm{r})=\sum_{\lambda}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\mathrm{in/out},\lambda}(\bm{r})a_{u}^{\mathrm{in/out},\lambda}(k)+\mathrm{H.c..}$ (36)
We provide the fields inside the microring resonators for each case in Fig. 4, obtained from the Heisenberg equations for the specific asymptotic state [27]. Using $\bm{r}_{\perp}$ to denote a vector along the cross-section of the waveguide constructing the ring, and $\zeta$ to identify the azimuthal coordinate along the circumference of the ring (see Fig. 1a), the asymptotic-in fields for the actual waveguide (Fig. 4a) are
$\bm{\mathcal{D}}_{uk}^{\mathrm{in,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u-}^{\mathrm{ac}}(k)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (37)
$\bm{\mathcal{D}}_{uk}^{\mathrm{in,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(2)}_{u}(\bm{r}_{\perp})F_{u-}^{\mathrm{ac}(2)}(k)\left(1+\frac{i\gamma_{u}^{\mathrm{ac}(1)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(1)}}F_{u-}^{\mathrm{ac}(1)}(k)\right)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (38)
the asymptotic-out fields for the actual waveguide (Fig. 4b) are
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ac}(1)}(k)\left(1-\frac{i\gamma_{u}^{\mathrm{ac}(2)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(2)}}F_{u+}^{\mathrm{ac}(2)}(k)\right)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (39)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(2)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ac}(2)}(k)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (40)
the asymptotic-out fields for the first phantom waveguide (Fig. 4 c)) are
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(1)}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ph}(1)}(k)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (41)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(1)}(\bm{r}_{\perp})=0,$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (42)
and the asymptotic-out fields for the second phantom waveguide (Fig. 4d) are
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(2)}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ac}(1)}(k)\left(-\frac{i\gamma_{u}^{\mathrm{ac}(2)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(2)}}F_{u+}^{\mathrm{ph}(2)}(k)\right)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta}$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (43)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(2)}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(2)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ph}(2)}(k)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (44)
where $\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r})$ are the spatial mode profiles in the $\ell^{\mathrm{th}}$ ring ($\ell\in\{1,2\}$) associated to the frequency range $u$ [24]. We have also introduced a field enhancement factor
${F}_{u\pm}^{\lambda(\ell)}(k)=\frac{1}{\sqrt{\mathcal{L}^{(\ell)}}}\left(\frac{({\gamma}_{u}^{\lambda(\ell)})^{*}}{v_{u}(K_{u}-k)\pm i\bar{\Gamma}_{u}^{(\ell)}}\right).$ (45)
The spatial modes are normalized according to
$\begin{split}\int\frac{v_{p}(\bm{r}_{\perp};\omega_{u})}{v_{g}(\bm{r}_{\perp};\omega_{u})}\frac{\bm{\mathsf{d}}_{u}^{(\ell)*}(\bm{r}_{\perp})\cdot\bm{\mathsf{d}}_{u}^{(\ell)}(\bm{r}_{\perp})}{\epsilon_{0}\varepsilon_{1}(\bm{r}_{\perp};\omega_{u})}\mathrm{d}\bm{r}_{\perp}=1,\end{split}$ (46)
where $v_{p}$ and $v_{g}$ are the local phase and group velocities of the material [24].
Figure 4: Schematic of the asymptotic-in state for the actual waveguide (a), the asymptotic-out for the actual waveguide (b), the asymptotic-out for the first phantom waveguide (c), and the asymptotic-out for the second phantom waveguide (d).
Appendix D Supplemental material 2: Tensor and nonlinear parameter
The III-V semiconductors of interest such as AlGaAs and InGaP belong to the $\bar{4}3$m point group and have a cubic symmetry [51]. In general, the $\chi^{(3)}$ tensor has 21 nonzero components, of which only one component is independent assuming Kleinman symmetry. We have $\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}=\chi^{(3)}_{y^{\prime}y^{\prime}y^{\prime}y^{\prime}}=\chi^{(3)}_{z^{\prime}z^{\prime}z^{\prime}z^{\prime}}=3\chi^{(3)}_{[x^{\prime}x^{\prime}y^{\prime}y^{\prime}]}=3\chi^{(3)}_{[x^{\prime}x^{\prime}z^{\prime}z^{\prime}]}=3\chi^{(3)}_{[y^{\prime}y^{\prime}z^{\prime}z^{\prime}]}=\bar{\chi}^{(3)}$, where the notation $[abcd]$ denotes the distinct permutations of the set $\{a,b,c,d\}$.
We can introduce a laboratory frame that coincides with the crystal’s axes $x^{\prime}y^{\prime}z^{\prime}$, and consider a crystal grown in the $z^{\prime}$ direction. We also introduce a ring frame xyzxyz where the $zx$ plane forms the cross-section of the waveguide constructing the ring, and $y$ is in the direction of the azimuthal angle going around the ring’s circumference (see Fig. 1). Since we assume Kleinman symmetry with a cubic crystal, the $\chi^{(3)}$ tensor takes a form similar to one of an isotropic material, and is therefore invariant under rotations of the coordinate system (e.g., about the $z^{\prime}$-axis). Consequently, in our case with the defined coordinate systems, the transformed tensor components satisfy $\chi^{(3)}_{i^{\prime}j^{\prime}k^{\prime}l^{\prime}}=\chi^{(3)}_{ijkl}$.
We can write Eqs. (28) and (29) in terms of nonlinear parameters $\gamma_{\mathrm{NL}}$ [24]. For simplicity, we can write the fields from Eqs. (37)–(44) as
$\bm{\mathcal{D}}_{uk}^{\mathrm{in/out},\lambda}(\bm{r})=\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r})\tilde{{\mathcal{D}}}_{uk}^{\mathrm{in/out},\lambda}e^{i\kappa^{(\ell)}_{u}\zeta}\,\,\,\,[\bm{r}\in\,\mathrm{ring}\,\ell],$ (47)
and we can write
$\begin{split}K_{1}^{\lambda\lambda^{\prime}}(k_{1},...,k_{4})&=\frac{\hbar^{2}\sqrt{\omega_{S_{1}}\omega_{I}}v_{P_{1}}v_{P_{1}}\mathcal{L}^{(1)}\gamma_{\mathrm{NL}}^{(1)}}{4\pi^{2}}\left[\tilde{{\mathcal{D}}}_{S_{1},k_{1}}^{\mathrm{out},\lambda}\tilde{{\mathcal{D}}}_{I,k_{2}}^{\mathrm{out},\lambda^{\prime}}\right]^{*}\tilde{{\mathcal{D}}}_{P_{1},k_{3}}^{\mathrm{in,ac}}\tilde{{\mathcal{D}}}_{P_{1},k_{4}}^{\mathrm{in,ac}}\end{split}$ (48)
$\begin{split}K_{2}^{\mu\mu^{\prime}\mu^{\prime\prime}}(k_{1},...,k_{4})&=\frac{\hbar^{2}\sqrt{\omega_{S_{2}}\omega_{S_{3}}}v_{I}v_{P_{2}}\mathcal{L}^{(2)}\gamma_{\mathrm{NL}}^{(2)}}{2\pi^{2}}\left[\tilde{{\mathcal{D}}}_{S_{2},k_{1}}^{\mathrm{out},\mu}\tilde{{\mathcal{D}}}_{S_{3},k_{2}}^{\mathrm{out},\mu^{\prime}}\right]^{*}\tilde{{\mathcal{D}}}_{I,k_{3}}^{\mathrm{out},\mu^{\prime\prime}}\tilde{{\mathcal{D}}}_{P_{2},k_{4}}^{\mathrm{in,ac}},\end{split}$ (49)
where we define
$\gamma_{\mathrm{NL}}^{(1)}$ $\equiv\frac{3\sqrt{\omega_{P_{1}}\omega_{P_{1}}}}{4\epsilon_{0}v_{P_{1}}v_{P_{1}}\mathcal{L}^{(1)}}\int\mathrm{d}\bm{r}\Gamma_{ijkl}^{(3)}(\bm{r})\left[{\mathsf{d}}^{(1),i}_{S_{1}}(\bm{r}){\mathsf{d}}^{(1),j}_{I}(\bm{r})\right]^{*}{\mathsf{d}}^{(1),k}_{P_{1}}(\bm{r}){\mathsf{d}}^{(1),l}_{P_{1}}(\bm{r})e^{i(\kappa^{(1)}_{P_{1}}+\kappa^{(1)}_{P_{1}}-\kappa^{(1)}_{I}-\kappa^{(1)}_{S_{1}})\zeta},$ (50)
$\gamma_{\mathrm{NL}}^{(2)}$ $\equiv\frac{3\sqrt{\omega_{I}\omega_{P_{2}}}}{4\epsilon_{0}v_{I}v_{P_{2}}\mathcal{L}^{(2)}}\int\mathrm{d}\bm{r}\Gamma_{ijkl}^{(3)}(\bm{r})\left[{\mathsf{d}}^{(2),i}_{S_{2}}(\bm{r}){\mathsf{d}}^{(2),j}_{S_{3}}(\bm{r})\right]^{*}{\mathsf{d}}^{(2),k}_{I}(\bm{r}){\mathsf{d}}^{(2),l}_{P_{2}}(\bm{r})e^{i(\kappa^{(2)}_{I}+\kappa^{(2)}_{P_{2}}-\kappa^{(2)}_{S_{3}}-\kappa^{(2)}_{S_{2}})\zeta}.$ (51)
To compute the nonlinear parameters, we relate the $\Gamma^{(3)}$ tensor to the nonlinear susceptibility tensor $\chi^{(3)}$
$\Gamma^{(3)}_{ijkl}(\bm{r})=\frac{\chi^{(3)}_{ijkl}(\bm{r})}{\epsilon_{0}^{2}\varepsilon_{1}^{4}(\bm{r})},$ (52)
where in identifying the effective nonlinear coefficient we neglect any frequency dependence on the dielectric constants, since we consider frequencies in the same range [24]. We can relate the displacement fields to the electric fields
$\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r})=\epsilon_{0}\varepsilon_{1}(\bm{r})\bm{\mathsf{e}}^{(\ell)}_{u}(\bm{r}_{\perp},\zeta),$ (53)
and we can write
$\begin{split}\gamma_{\mathrm{NL}}^{(\ell)}&=\frac{3\bar{\chi}^{(3)}\epsilon_{0}\sqrt{\omega_{{u_{3}}}\omega_{{u_{4}}}}}{4v_{{u_{3}}}v_{{u_{4}}}\mathcal{L}^{(\ell)}}\int\mathrm{d}\bm{r}_{\perp}\mathrm{d}\zeta\frac{\chi^{(3)}_{ijkl}(\bm{r})}{\bar{\chi}^{(3)}}\left[{\mathsf{e}}^{(\ell),i}_{{u_{1}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),j}_{{u_{2}}}(\bm{r}_{\perp})\right]^{*}{\mathsf{e}}^{(\ell),k}_{{u_{3}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),l}_{{u_{4}}}(\bm{r}_{\perp})e^{i(\kappa^{(\ell)}_{{u_{3}}}+\kappa^{(\ell)}_{{u_{4}}}-\kappa^{(\ell)}_{{u_{2}}}-\kappa^{(\ell)}_{{u_{1}}})\zeta},\end{split}$ (54)
where $\bm{{\mathsf{e}}}^{(\ell)}_{u}(\bm{r}_{\perp})$ can be extracted from an eigenmode solver for the specific structure and the integral over $\bm{r}_{\perp}$ can be evaluated numerically. We have used the fact that in the ring frame we can write the mode profiles as $\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r}_{\perp})$, since we assume the spatial profiles are independent of $\zeta$. For phase matched wavenumbers $\kappa^{(\ell)}_{{u_{3}}}+\kappa^{(\ell)}_{{u_{4}}}-\kappa^{(\ell)}_{{u_{2}}}-\kappa^{(\ell)}_{{u_{1}}}=0$, and we have
$\begin{split}\gamma_{\mathrm{NL}}^{(\ell)}&=\frac{3\bar{\chi}^{(3)}\epsilon_{0}\sqrt{\omega_{{u_{3}}}\omega_{{u_{4}}}}}{4v_{{u_{3}}}v_{{u_{4}}}}\int\mathrm{d}\bm{r}_{\perp}\frac{\chi^{(3)}_{ijkl}(\bm{r})}{\bar{\chi}^{(3)}}\left[{\mathsf{e}}^{(\ell),i}_{{u_{1}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),j}_{{u_{2}}}(\bm{r}_{\perp})\right]^{*}{\mathsf{e}}^{(\ell),k}_{{u_{3}}}(\bm{r}_{\perp}){\mathsf{e}}^{(\ell),l}_{{u_{4}}}(\bm{r}_{\perp}).\end{split}$ (55)